# Stacked Intelligent Metasurface for End-to-End OFDM System

### v1

@@ III E2E problem-solving
$\mathbf{\Theta}_{p}^{l}\leftarrow\mathbf{\Theta}_{p}^{l}-\eta\nabla_{\mathbf{\Theta}_{p}^{l}}\mathcal{L}\left(\mathbf{b},\hat{\mathbf{b}}\right),p\in\left\{0,1\right\},1\leqslant l\leqslant L,$ (36)
$\mathbf{\Xi}_{p}^{k}\leftarrow\mathbf{\Xi}_{p}^{k}-\eta\nabla_{\mathbf{\Xi}_{p}^{k}}\mathcal{L}\left(\mathbf{b},\hat{\mathbf{b}}\right),p\in\left\{0,1\right\},1\leqslant k\leqslant K,$ (37)
@@
As shown in Fig. 7, we propose a training and deployment framework for the EMNN. Importantly, the EMNN is always trained on the BS side. At regular intervals, the BS prunes and quantizes the weights of the RX-SIM (RX-DPSIM) NN and the UE-DNN, and then distributes them to the corresponding users. Finally, the BS transmitter deploys the BS-DNN in real time and dynamically controls the EM (DPEM) units according to the weights of the TX-SIM (TX-DPSIM) NN. Each user, in turn, deploys the UE-DNN based on the received weights and simultaneously adjusts the EM (DPEM) units in the RX-SIM (RX-DPSIM) dynamically. The specific training and deployment steps of the EMNN are as follows
@@
Step 2. Instantaneous CSI acquisition and online training of EMNN. The BS obtains the instantaneous CSI of each user through channel estimation, channel prediction, and other methods. Then, the EMNN performs real-time online training through the instantaneous channel layer.
Step 3. Parameter distribution. At regular short intervals, the BS extracts and partitions the weights of the EMNN, retaining those corresponding to the transmitter, while distributing the weights associated with the receiver to the respective users.
Step 4. Systems control. BS-DNN and UE-DNN are conventional neural networks that can be directly deployed inside the transceivers based on the distributed weight data, whereas the TX-SIM (TX-DPSIM) and RX-SIM (RX-DPSIM) devices require solving the weight data according to predefined mapping rules before enabling real-time control of the EM (DPEM) units.
@@
The overall training complexity of the proposed EMNN arises from the superposition of wave-domain signal processing in the metasurface network (SIM-NN/DPSIM-NN) and digital computation in conventional neural networks (BS-DNN/UE-DNN). For the SIM-assisted system, the complexity can be decomposed into several components. The complexity of the BS-DNN and UE-DNN depends on their network architectures. If the typical hidden layer dimension is denoted by $D$, the number of parameters is on the order of $\mathcal{O}(D^{2}HJ)$, where $H$ denotes the number of hidden layers and $J$ denotes the number of users. The computational complexity of forward propagation is mainly dominated by matrix multiplications and is also on the order of $\mathcal{O}(D^{2}HJ)$. For the metasurface part, the number of parameters is $\mathcal{O}(ML)$ for the TX-SIM and $\mathcal{O}(NKJ)$ for all RX-SIMs at the users. The corresponding computational complexity is dominated by matrix multiplications simulating the electromagnetic transmission between analog layers. For each subcarrier, the complexity of the TX-SIM is on the order of $\mathcal{O}(M^{2}L)$, while that of all RX-SIMs is on the order of $\mathcal{O}(N^{2}KJ)$. Consequently, the total number of parameters of the SIM-assisted system is $\mathcal{O}(ML+NKJ+D^{2}HJ)$, and the overall computational complexity of a single forward propagation, scaling with the number of subcarriers $N^{c}$, is given by $\mathcal{O}(N^{c}(M^{2}L+N^{2}KJ)+D^{2}HJ)$.
Similarly, for the DPSIM-assisted system, its architecture can be regarded as two parallel SIMs, resulting in most of its structural complexities being approximately twice those of the SIM-assisted system. Assuming that the DNN architectures in the DPSIM-assisted system are identical to those in the SIM-assisted system, the total number of parameters for the DPSIM-assisted system is $\mathcal{O}\!\left(2(M^{\mathrm{dp}}L^{\mathrm{dp}}+N^{\mathrm{dp}}K^{\mathrm{dp}}J)+D^{2}HJ\right)$, and the overall computational complexity is approximately $\mathcal{O}\!\left(2N^{c}\left((M^{\mathrm{dp}})^{2}L^{\mathrm{dp}}+(N^{\mathrm{dp}})^{2}K^{\mathrm{dp}}J\right)+D^{2}HJ\right)$.
@@
Considering that the framework proposed in Fig. 7 requires sophisticated cross-layer modeling and system-level implementation, this section focuses on verifying its core component, namely the EMNN-based transfer learning approach, in both SIM-assisted E2E OFDM systems and DPSIM-assisted E2E OFDM systems. The performance of this method plays a decisive role in the overall effectiveness of the proposed framework, and thus its validation results largely reflect the feasibility of the framework. Specifically, we assume a quasi-static block fading channel. First, the EMNN is trained as a baseline model using the statistical channel. Then, several sets of instantaneous channels are randomly generated, where the channel layer in the EMNN is accordingly replaced, and the model is fine-tuned via transfer learning. Finally, the average testing performance under different channel conditions is adopted as the system performance metric.
Figure 8: Performance comparison chart of different training methods.
@@
where $\text{PL}(d_{0})=20\log_{10}\left(\frac{4\pi d_{0}}{\lambda}\right)\text{ dB}$ is the free space path loss at the reference distance $d_{0}$, $b$ represents the path loss exponent, $X_{\delta}$ is a zero mean Gaussian random variable with a standard deviation $\delta$, characterizing the large-scale signal fluctuations of shadow fading. Furthermore, Table II sequentially summarizes the system parameter settings, channel parameter settings, SIM (DPSIM) parameter settings, and model training/testing parameter settings used for simulation. Unless otherwise specified, the subsequent simulation results are obtained based on the above settings.
@@
Figure 10: Performance comparison chart of different EM (DPEM) unit quantities.
Figure 11: Performance comparison chart of different metasurface layer unit quantities.
@@
Fig. 10 – Fig. 12 depict the BER curves with varying numbers of EM (DPEM) units, metasurface layers, and transceiver antennas, respectively. It can be observed that enlarging the EM (DPEM) array size, increasing the number of metasurface layers, and employing more transmit/receive antennas all contribute to effectively reducing the system BER. Moreover, under all configurations, the DPSIM-assisted E2E OFDM system consistently achieves significantly better BER performance than its SIM-assisted counterpart. Particularly, we observe that by integrating the SIM (DPSIM) device, an E2E OFDM system with 16T9R can achieve comparable performance to a massive MIMO OFDM system with 256T49R. This validates the potential of incorporating SIM (DPSIM) devices for wave-domain signal processing to offload baseband processing tasks, thereby simplifying the transceiver design and reducing hardware cost.
Figure 12: Performance comparison chart of different antenna quantities.
Figure 13: Performance comparison chart of different subcarrier quantities. Massive MIMO (256T49R).
@@
Figure 14: Performance comparison chart of different transmission bit numbers. Massive MIMO (256T49R).
Figure 15: Performance comparison chart of different polarization radiation factors. Massive MIMO (256T49R).
V Conclusion

### v2

@@ Abstract
$\mathbf{\Theta}_{p}^{l}\leftarrow\mathbf{\Theta}_{p}^{l}-\eta\nabla_{\mathbf{\Theta}_{p}^{l}}\mathcal{L}\left(\mathbf{b},\hat{\mathbf{b}}\right),1\leqslant l\leqslant L^{\mathrm{dp}},$ (36)
$\mathbf{\Xi}_{p}^{k}\leftarrow\mathbf{\Xi}_{p}^{k}-\eta\nabla_{\mathbf{\Xi}_{p}^{k}}\mathcal{L}\left(\mathbf{b},\hat{\mathbf{b}}\right),1\leqslant k\leqslant K^{\mathrm{dp}},$ (37)
@@
As shown in Fig. 7, we propose a training and deployment framework for EMNN. Importantly, EMNN is always trained on the BS side. At regular intervals, the BS prunes and quantizes the weights of the RX-SIM (RX-DPSIM) NN and the UE-DNN, and then distributes them to the corresponding users. Finally, the BS transmitter deploys the BS-DNN in real time and dynamically controls the EM (DPEM) units according to the weights of the TX-SIM (TX-DPSIM) NN. Each user, upon receiving the weights, deploys the UE-DNN and simultaneously adjusts the EM (DPEM) units in the RX-SIM (RX-DPSIM) dynamically. The specific training and deployment procedures of EMNN are as follows
@@
Step 2. Instantaneous CSI acquisition and online training of EMNN. The BS obtains the instantaneous CSI of each user through channel estimation or channel prediction. Then, the EMNN performs real-time online training through the instantaneous channel layer.
Step 3. Parameter distribution. At short regular intervals, the BS extracts and partitions the EMNN weights, retaining the weights corresponding to the transmitter side, while distributing the receiver-side weights to the respective users.
Step 4. Systems control. BS-DNN and UE-DNN are conventional neural networks that can be directly deployed inside the transceivers based on the distributed weight data, whereas the TX-SIM (TX-DPSIM) and RX-SIM (RX-DPSIM) devices must apply an inverse transformation to the weight data according to predefined mapping rules before real-time control of the EM (DPEM) units.
@@
The overall training complexity of the proposed EMNN arises from the superposition of wave-domain signal processing in the metasurface network and digital computation in conventional neural networks. For the SIM-assisted system, the overall complexity can be decomposed into several components. In particular, the complexities of the BS-DNN and UE-DNN are primarily determined by their respective network architectures. If the typical hidden layer dimension is denoted by $D$, the number of parameters is on the order of $\mathcal{O}(D^{2}HJ)$, where $H$ denotes the number of hidden layers. The computational complexity of forward propagation is mainly dominated by matrix multiplications and is also on the order of $\mathcal{O}(D^{2}HJ)$. For the metasurface part, the number of parameters is $\mathcal{O}(ML)$ for the TX-SIM and $\mathcal{O}(NKJ)$ for all RX-SIMs at the users. The corresponding computational complexity is dominated by matrix multiplications simulating the EM transmission between analog layers. For each subcarrier, the complexity of the TX-SIM is on the order of $\mathcal{O}(M^{2}L)$, while that of all RX-SIMs is on the order of $\mathcal{O}(N^{2}KJ)$. Consequently, the total number of parameters of the SIM-assisted system is $\mathcal{O}(ML+NKJ+D^{2}HJ)$, and the overall computational complexity of a single forward propagation, scaling with the number of subcarriers $N^{\mathrm{c}}$, is given by $\mathcal{O}(N^{\mathrm{c}}(M^{2}L+N^{2}KJ)+D^{2}HJ)$.
Similarly, for the DPSIM-assisted system, its architecture can be regarded as two parallel SIMs, resulting in most of its structural complexities being approximately twice those of the SIM-assisted system. Assuming that the DNN architectures in the DPSIM-assisted system are identical to those in the SIM-assisted system, the total number of parameters for the DPSIM-assisted system is $\mathcal{O}\!\left(2(M^{\mathrm{dp}}L^{\mathrm{dp}}+N^{\mathrm{dp}}K^{\mathrm{dp}}J)+D^{2}HJ\right)$, and the overall computational complexity is approximately $\mathcal{O}\!\left(2N^{\mathrm{c}}(M^{\mathrm{dp}})^{2}L^{\mathrm{dp}}+2N^{\mathrm{c}}(N^{\mathrm{dp}})^{2}K^{\mathrm{dp}}J+D^{2}HJ\right)$.
@@
Considering that the framework proposed in Fig. 7 requires complex cross-layer modeling and system-level implementation, in this section we focus on validating its core component, the EMNN method trained via transfer learning in SIM-assisted E2E OFDM systems and DPSIM-assisted E2E OFDM systems. The performance of this method plays a decisive role in the overall effectiveness of the framework; therefore, its validation results largely reflect the feasibility of the proposed framework. Specifically, we assume a quasi-static block-fading channel. First, the EMNN uses statistical CSI to train a base model. Then, multiple sets of instantaneous channels are randomly generated, and for each set, the EMNN replaces the channel layer accordingly and fine-tunes the model via transfer learning. Finally, the average of the test results under different channel conditions is taken as the system performance metric.
Figure 8: Performance comparison of different training methods.
@@
where $\text{PL}(d_{0})=20\log_{10}\left(\frac{4\pi d_{0}}{\lambda}\right)\text{ dB}$ is the free space path loss at the reference distance $d_{0}$, $b$ represents the path loss exponent, $X_{\delta}$ is a zero mean Gaussian random variable with a standard deviation $\delta$, characterizing the large-scale signal fluctuations of shadow fading. Furthermore, Table II sequentially summarizes the system parameter settings, channel parameter settings, SIM (DPSIM) parameter settings, and model training (testing) parameter settings used for simulation. Unless otherwise specified, the subsequent simulation results are obtained based on the above settings.
@@
Figure 10: Performance comparison of different EM (DPEM) unit quantities.
Figure 11: Performance comparison of different numbers of metasurface layers.
@@
Fig. 10 – Fig. 12 depict the BER curves with varying numbers of EM (DPEM) units, metasurface layers, and transceiver antennas, respectively. It can be observed that enlarging the EM (DPEM) array size, increasing the number of metasurface layers, and employing more transmit (receive) antennas all contribute to effectively reducing the system BER. Moreover, under all configurations, the DPSIM-assisted E2E OFDM system consistently achieves significantly better BER performance than its SIM-assisted counterpart. Particularly, we observe that by integrating the SIM (DPSIM) device, an E2E OFDM system with 16T9R can achieve comparable performance to a massive MIMO OFDM system with 256T49R. This validates the potential of incorporating SIM (DPSIM) devices for wave-domain signal processing to offload baseband processing tasks, thereby simplifying the transceiver design and reducing hardware cost.
Figure 12: Performance comparison of different antenna quantities.
Figure 13: Performance comparison of different subcarrier quantities. Massive MIMO (256T49R).
@@
Figure 14: Performance comparison of different transmission bit numbers. Massive MIMO (256T49R).
Figure 15: Performance comparison of different polarization radiation factors. Massive MIMO (256T49R).
V Conclusion.